A. Elimination Method; 1. 2x-5y = 9; 3x+4y = 6 2. 3x+4y= 3; 4x-5y=5


1
Expert's answer
2022-03-14T19:37:22-0400

"1. \\
2x-5y=9\\
3x+4y=6\\
\left \{\begin{matrix}
 4(2x-5y)=4\cdot9 \\
 5(3x+4y)=5\cdot6
\end{matrix}\right.\\
\left \{\begin{matrix}
 8x-20y=36 \\
 15x+20y=30
\end{matrix}\right.\\
23x=66\\
x=\frac{66}{23}\\
2\cdot\frac{66}{23}-5y=9\\
5y=\frac{132}{23}-9\\
5y=-\frac{75}{23}\\
y=-\frac{15}{23}\\(\frac{66}{23},-\frac{15}{23})"

"2. \\
3x+4y=3\\
4x-5y=5\\
\left \{\begin{matrix}
 5(3x+4y)=5\cdot3 \\
 4(4x-5y)=4\cdot5
\end{matrix}\right.\\
\left \{\begin{matrix}
 15x+20y=15 \\
 16x-20y=20
\end{matrix}\right.\\
31x=35\\
x=\frac{35}{31}\\
3\cdot\frac{35}{31}+4y=3\\
4y=3-\frac{105}{31}\\
4y=-\frac{12}{31}\\
y=-\frac{3}{31}\\(\frac{35}{31},-\frac{3}{31})"